DO NOT USE EXCEL. APPLY FINANCIAL MATH CONCEPT (BOND):

The net present values of the following two projects are equal at an annual effective interest rate of 8%. Project A requires an investment of $5000 today and pays 2000 in one year and 5000 in five years. Project B requires an investment of X today and pays 3000 in two years and 7000 in eight years. Calculate X.
Answer #1

Solution :

It is given that NPV of both the projects are same hence we need to find the NPV of A and B and equate it to find the value of X.

Lets find nipv of both the projects NPV of Project A NPVA = -5000 + 2000 + 5000 [1+8.1)* ( 1+80) 5 = -5000 + 2008 + soos :

Value of X is = $6099.13